OpenStudy (anonymous):

what is the cube root of 56x to the fifth power times y.

OpenStudy (anonymous):

\[\sqrt[3]{56}=\sqrt[3]{8\times 7}=\sqrt[3]{8}\times \sqrt[3]{7}=2\sqrt[3]{7}\] is as start

OpenStudy (anonymous):

also \[\sqrt[3]{x^5}=\sqrt[3]{x^3x^2}=x\sqrt[3]{x^2}\]

OpenStudy (anonymous):

so a two comes out, an x comes out, and there is some stuff left in. final answer is \[2x\sqrt[3]{x^2y}\]

OpenStudy (anonymous):

you did cube root of 56 but where is to the fifth power times y? :(